1. ***present all the calculations***

Part I: The density of seawater is approximately 1.027g / cm3 and the ice density 0.93g / cm3. United Nations
Iceberg (iceberg), generally has a mass of 150,000 metric tons (1 metric ton = 1000kg). Calculation
(a) The buoyant force exerted by the water on the "iceberg",
(b) the volume of the iceberg above sea level,
(c) the volume fraction above sea level.

Part II:

A metal bucket with a height of 0.700 m and a mass of 650 kg is suspended from a rope in an open tank containing a liquid with a density equal to 1530 kg / m3. The block is submerged 0.300 m below the surface. Calculate:
(a) The force up at the bottom of the block,
(b) the tension force of the rope,
(c) the magnitude of the buoyant force in the cube using the Archimedes principle.
Answer #1

Part I)

ρwater=1027kg/m3, ρice=930kg/m3

a)

Vtotal = miceice =(150,000,000kg)/(930kg/m3)=161290.3 m3

Fo find buoyant force use equation,

Fb = Fg

Fb = ρiceVtotalg

Plug values,

Fb = (930kg/m3)*(161290.3m3)*(9.8m/s2)

Fb = 1.47*10^9 N

b)

Use equation,

Fb = Fg

ρwaterVbelowg= ρiceVtotalg

ρwaterVbelow= ρiceVtotal

Vbelow=Vtotalicewater) -------------(1)

Plug values,

Vbelow= (161290.3 m3)*(930kg/m3/1027kg/m3)

Vbelow = 146056.45 m3

Vabove = Vtotal - Vbelow = 161290.3 m3 - 146056.45 m3 = 15233.85 m3

c)

From eqn (1),

Vbelow/Vtotalicewater

Vbelow/Vtotal=(930kg/m3/1027kg/m3)

Vbelow/Vtotal= 0.9055

Vabove/Vtotal= 1- Vbelow/Vtotal

Vabove/Vtotal = 1- 0.9055 = 0.0945

In percentage it is 9.45%